# A framework for the valuation of insurance liabilities by production cost

### v1

@@ Abstract
This paper sets out a framework for the valuation of insurance liabilities that is intended to be economically realistic, elementary, reasonably practically applicable, and as a special case to provide a basis for the valuation in regulatory solvency systems such as Solvency II and the SST. The valuation framework is based on the cost of producing the liabilities to an insurance company that is subject to solvency regulation (regulatory solvency capital requirements) and insolvency laws (consequences of failure) in finite discrete time. Starting from the replication approach of classical no-arbitrage theory, the framework additionally considers the nature and cost of capital (expressed by a “financiability condition”), that the liabilities may be required to be fulfilled only “in sufficiently many cases” (expressed by a “fulfillment condition”), production using “fully illiquid” assets in addition to tradables, and the asymmetry between assets and liabilities. We identify necessary and sufficient conditions on the capital investment under which the framework recovers the market prices of tradables, investigate extending production to take account of insolvency, implications of using illiquid assets in the production, and show how Solvency II and SST valuation can be derived with specific assumptions.
@@ 1 Introduction
Insurance contracts promise the payment of future claims and benefits whose amounts are characteristically currently uncertain, depending for example on whether a person dies, whether an earthquake occurs and the resulting damage, or on the development of financial markets. A key task specifically for regulatory solvency capital requirements such as Solvency II ([3], [6]) and the Swiss Solvency Test (SST, [9]) is to assess whether an insurance company is able to fulfill its promises made under insurance contracts. Importantly, insurance companies are in general not able to fulfill their insurance obligations with certainty, and regulatory solvency capital requirements such as in Solvency II and the SST can be interpreted as requiring fulfillment roughly “in sufficiently many cases”, expressed by what we call “fulfillment condition” in this paper. The fulfillment condition in Solvency II is given by a 99.5% value-at-risk and in the SST by a 99% tail-value-at-risk (expected shortfall), in both cases over a one-year period. This also means that insurance companies may fail, i.e. default on their obligations or become balance sheet insolvent, which triggers insolvency laws, with potentially serious consequences such as bankruptcy liquidation.
To assess whether an insurance company is able to fulfill its insurance obligations in sufficiently many cases, it is natural to try to find suitable “production strategies” with which the insurance company can produce the insurance obligations in sufficiently many cases. A (financial) production strategy roughly consists in dynamically buying and selling assets (and possibly liabilities), including outstanding incoming premiums, and fulfilling insurance obligations by selling assets or paying out incoming payouts from assets such as bond coupons or dividends. When an insurance company can finance the production cost for a suitable production strategy, it can fulfill the insurance obligations in sufficiently many cases using this strategy. The value of the insurance liabilities can be defined as the minimum (or infimum) of the production cost over suitable production strategies. This idea arguably underlies replication or hedging strategies considered in classical no-arbitrage theory.
This paper attempts to set out a reasonably practically applicable and elementary framework for the valuation of insurance liabilities that is based on production cost for an insurance company in an economically realistic setting, specifically taking account of solvency regulation (regulatory capital requirements) and insolvency laws (consequences of failure), and providing a basis for valuation approaches in regulatory solvency systems such as Solvency II and the SST. It takes place in finite discrete time and starts from known stochastic insurance liability cash flows.
@@
The range of possible production strategies is reduced because typically not all assets held by insurance companies are tradables, that is, “fully liquid”. Illiquid assets, including recoverables from reinsurance, have limited tradability, so they can potentially not be exchanged for other assets or sold for paying insurance obligations. In this paper, we assume that the assets of insurance companies consist of tradables and “fully illiquid assets” that we assume can neither be bought nor sold.111“Fully illiquid assets” in this sense are linked to discussions in the insurance industry about whether insurance companies can benefit from “holding assets to maturity” (HTM) instead of having to sell them for a potentially unfavorable market price, which under Solvency II is related to “volatility adjustment” and “matching adjustment”.
The range of possible production strategies is significantly extended by taking into account that insurance companies have and use capital, e.g. shareholder capital. Capital (more generally hybrid capital) can be seen as non-binding promises of payouts to the capital investors (for hybrid capital, non-binding under specific circumstances) in the sense that non-payment does not trigger insolvency proceedings. Capital can thus contribute to production because payouts can be reduced as required in any (or some) circumstances and instead be used for producing the insurance obligations. In return, the stochastic future payouts to the capital investors have to be sufficient overall, which we specify by what we call “financiability condition”. In line with the conception and the one-year view of Solvency II and the SST, we make the strong assumption that capital can be raised annually if required provided that the financiability condition is fulfilled.
Our framework starts from the replication approach of classical no-arbitrage theory and extends it to consider the fulfillment of obligations only in sufficiently many cases, the nature and cost of capital, and fully illiquid assets in addition to tradables. The framework also considers insolvency, specifically that companies can default on obligations. This creates an asymmetry between assets and liabilities: going short in a tradable asset as required by typical no-arbitrage arguments in reality means taking on a liability that needs to be precisely defined and is in important ways different from a long position in the asset; specifically, it can trigger insolvency proceedings. Similarly, there is a difference between contractual cash flows on the one hand and actual cash flows that account for counterparty default.
One of the main results of this paper (Theorem 5.3) characterizes the conditions on the capital investment under which the valuation according to the framework recovers the market prices of tradables, more precisely, the market values of non-negative investment strategies of tradables. For this setting, fulfillment of obligations only in sufficiently many cases and illiquid assets are not considered and it is assumed that short positions in tradables as liabilities can be taken on and closed out for the market price at any point in time.
A key element of the framework is the precise definition of production strategies. We assume that production strategies proceed over successive one-year periods, in line with the one-year view of regulatory capital requirements such as Solvency II and the SST. Sightly simplified, under a production strategy for a given fulfillment and financiability condition, at the start $i$ of a one-year period from date $i$ to $i+1$, the company has assets with value equal to or exceeding the production cost of the insurance liabilities, with the excess representing capital. Under the strategy, from date $i$ to $i+1$, assets are traded and used to pay obligations by selling assets (for tradables only) or paying out incoming cash flows from the assets (also for fully illiquid assets). By construction of the strategy, at the end $i+1$ of the one-year period, in sufficiently many cases as defined by the fulfillment condition, the value of the assets is at least as large as the remaining production cost of the liabilities. Any positive excess represents the claim of the capital investors, where by construction, seen from date $i$, the stochastic excess satisfies the financiability condition. By construction, at $i+1$, if needed, additional capital in the form of additional tradable assets can be raised from capital investors up to the level required for the fulfillment condition to be satisfied at the end $i+2$ of the next period from date $i+1$ to $i+2$. In the remaining cases at date $i+1$, the excess is negative, so capital investors get nothing and the company fails through balance sheet insolvency, triggering insolvency laws. We consider the case that the insolvency laws specify proportionally reducing the outstanding liabilities so that these adjusted liabilities can be fulfilled in sufficiently many cases. For insureds, this may arguably be preferable to bankruptcy liquidation. We find that, for a positively homogeneous financiability condition, the suitably adjusted production strategy allows producing the adjusted liabilities in all cases.
The construction of production strategies proceeds recursively backward in time, starting at the end of the lifetime of the insurance liabilities. The basic idea consists of the following two steps. By recursion, the (stochastic) production cost at the end $i+1$ of a one-year period from date $i$ to $i+1$ are known. In a first step, a production strategy over the one-year period needs to be constructed such that, at $i+1$, the fulfillment condition is satisfied, i.e. in sufficiently many cases, the value of the assets is larger than or equal to the known production cost at $i+1$. In a second step, the assets from the production strategy at the start $i$ of the one-year period are split into a (maximal) proportion that can be raised from capital investors and the remainder. The proportion that can be raised from capital investors is determined by the financiability condition and corresponds to the value to the capital investors at date $i$ of the stochastic claim to capital investors at $i+1$, given by the positive part of the excess of the value of the assets over the production cost at $i+1$. This works because the stochastic claim is independent of the proportion raised from capital investors. The remainder corresponds to the production cost at date $i$, allowing the recursion to proceed. The stronger the financiability condition (i.e. the higher the cost of capital), the higher the production cost, because a smaller fraction of the assets required at date $i$ can be raised for the given payout at $i+1$.
The framework of this paper shares basic ideas with but attempts to improve and generalize the earlier paper Möhr [8]. A setting similar to the earlier paper is used in Section 6, where we consider Solvency II and the SST as special cases. We refer to the introduction section of Engsner et al. [4] and the references provided there for some of the many papers that exist on the valuation of (insurance) liabilities. As in [4], the current paper considers successive one-year capital requirements and insolvency and, in common also with Albrecher et al. [2], a key element is the definition and valuation of capital investments.
In the current paper, in line with Möhr [8] and unlike several other approaches to the valuation of insurance liabilities, the value of the liabilities is considered as a whole. A breakdown into e.g. a best estimate and a risk margin is not an essential element. A key difference to [8], as well as to Engsner et al. [4] and others including Albrecher et al. [2], is that the production strategies are not restricted by assuming that the assets corresponding to the capital are invested separately and specifically in a numéraire asset. The current paper is intentionally more elementary than many other papers, not assuming the existence of (“risk-neutral”) pricing measures and numéraires. A specific difference to Engsner et al. [4] is the specification of the capital investment, which in our case is set up for every one-year period with payout equal to the positive part of the value of assets over liabilities. In [4], it consists of dividend streams until the option to default is exercised by capital investors and, in our understanding, does not consider the liability cash flows after the date considered.
The paper is organized as follows. In Section 3, we set up the framework, specifically cash flows, tradables, illiquid assets, and investment strategies, with specific consideration of portfolios containing negative units of tradables. We introduce a normative condition roughly comparable to “no arbitrage”, which we call “consistency”. Consistency essentially means that if, at a given point in time,, the market price of a suitable strategy is not smaller than that of another suitable strategy (almost surely), then this holds at any earlier point in time.
Section 4 sets out the production of liabilities in the framework, introducing the valuation of assets and liabilities in the balance sheet, fulfillment and the financiability condition, and the central definition of a production strategy, at first limited to portfolios with non-negative units of tradables and then in Section 4.5 extended to portfolios that can contain negative units. In Section 4.2, we show how, under suitable assumptions about the consequences of failures in insolvency laws, a production strategy for an arbitrary fulfillment condition can be extended to a production strategy covering all cases for an adjusted liability.
In Section 5, we analyze properties of the valuation, specifically, in Section 5.1, under which conditions the market prices of tradables are recovered. Section 5.2 looks at adding short positions to liabilities and Section 5.3 investigates production also using illiquid assets. In Section 6, we set out how the (capital cost) valuation of insurance liabilities under Solvency II and in the SST can be seen as special cases of the framework. Section 7 concludes.
2 Simple illustrative example
As a very simple illustration of basic ideas from Möhr [8] to which we come back in Section 6, consider the task of determining the value $v_{0}(L_{1})$ at date 00 of a given liability $L_{1}$ at date $1$ (with no cash flows between 00 and $1$) in terms of the cost of a strategy to produce $L_{1}$. Let $r_{0,1}$ be the “risk-free rate” at date 00 for a deterministic payoff at date $1$. Assume as fulfillment condition that the liabilities $L_{1}$ need to be fulfilled by the assets $A_{1}$ in the sense that $\rho[A_{1}-L_{1}]\leq 0$, where $\rho$ is a translation-invariant risk measure [^2] Translation invariance: $\rho[X+a]=\rho[X]-a$ for $a\in\mathbb{R}$., e.g. the $99.5\%$ value-at-risk. Let the strategy be defined by assets with value $A_{0}$ at date 00 invested risk-free, i.e. $A_{1}=(1+r_{0,1})\cdot A_{0}$. Then, the fulfillment condition is satisfied with equality for $A_{0}=(1+r_{0,1})^{-1}\cdot\rho[-L_{1}]$ by translation-invariance. Let the capital investors’ claim at date $1$ be $(A_{1}-L_{1})_{+}$ and assume that this quantity is positive with positive probability. Then, a capital investor would likely be prepared to invest a positive amount $C_{0}>0$ at date 00 for the claim $(A_{1}-L_{1})_{+}$ (financiability condition). We then set the production cost under the strategy as $v_{0}(L_{1})=A_{0}-C_{0}$.
@@ 2 Simple illustrative example
Further simplifying for illustration, let the financiability condition be given by an expected excess return $\eta>0$ over the risk-free rate and, even further simplifying, we ignore $()_{+}$, i.e. $E[A_{1}-L_{1}]=(1+r_{0,1}+\eta)\cdot C_{0}$. It then follows from a short calculation that the production cost under the strategy can be written as
$v_{0}(L_{1})=\frac{E[L_{1}]}{1+r_{0,1}}+\frac{\eta}{1+r_{0,1}+\eta}\cdot\rho\left[\frac{E[L_{1}]-L_{1}}{1+r_{0,1}}\right]$ (1)
@@
i.e. as the sum of a “best estimate” and a “risk margin” given by the discounted capital cost on the discounted deviation of $L_{1}$ from $E[L_{1}]$.

### v2

@@ Abstract
This paper sets out a framework for the valuation of liabilities from insurance contracts that is intended to be economically realistic, elementary, reasonably practically applicable, and as a special case to provide a basis for the valuation in regulatory solvency systems such as Solvency II and the SST. The valuation framework is based on the cost of producing the liabilities to an insurance company that is subject to solvency regulation (regulatory solvency capital requirements) and insolvency laws (definition and consequences of failure) in finite discrete time. Starting from the replication approach of classical no-arbitrage theory, the framework additionally considers the nature and cost of capital (expressed by a “financiability condition”), that the liabilities may only be required to be fulfilled “in sufficiently many cases” (expressed by a “fulfillment condition”), production using “fully illiquid” assets in addition to tradables, and the asymmetry between assets and liabilities. We identify the necessary and sufficient property of the financiability condition such that the framework recovers the market prices of tradables. We investigate extending production to take account of insolvency and implications of using illiquid assets in the production. We show how Solvency II and SST valuation can be derived with specific assumptions.
@@ 1 Introduction
Insurance contracts can be viewed as arrangements under which an insurance company promises the payment of future claims and benefits to policyholders and beneficiaries, for premiums paid by the policyholders. Characteristically, the amounts of future claims and benefits are currently uncertain, whereas the premiums are deterministic or at least “less uncertain”. Claims and benefits may for example depend on the damage caused by natural perils such as windstorms or floods, the death of a person, or the development of financial markets. Insurance contracts thus correspond to stochastic future cash flows, with cash inflows to the insurance company for premiums, and cash outflows for claims and benefits. In this paper, we assume that the stochastic cash flows are given and consider finite discrete time.
The objective of the paper is to set out an elementary and reasonably practically applicable framework for the valuation of insurance contracts in an economically realistic setting. All insurance contracts of an insurance company are valued jointly. The framework is intended as a special case to provide a basis for valuation approaches in regulatory solvency systems such as Solvency II ([3], [6]) and the Swiss Solvency Test (SST, [9]). In the framework, the value is based on the cost to the insurance company of fulfilling the promises made under the insurance contracts, i.e. of producing the insurance contract cash outflows (production cost). The insurance company produces the insurance contract cash outflows according to a production strategy. Under the production strategy, it dynamically trades assets and produces the insurance contract cash outflows by selling assets or paying out cash inflows such as premiums, bond coupons, or equity dividends.
This concept of production strategies is similar to replication or hedging strategies considered in classical no-arbitrage theory. To capture an economically realistic setting, we consider the following four additional aspects:
(1) Insurance companies may not be able to fulfill their insurance obligations, i.e. the promises made under the insurance contracts, with certainty, i.e. “in every possible state of the world”. For example, they may not be able to fulfill the maximum of all their claims and benefits. [^1] Diversification is a core element of the business model of insurance companies, which should make it unlikely that all future claims and benefits would simultaneously realize at their maximum. In turn, insurance companies are subject to solvency regulation and supervision and to insolvency laws. Insolvency laws determine when a company fails - typically, when it defaults on an obligation that is due or becomes balance sheet insolvent - and the consequences of failure, for example bankruptcy liquidation. Solvency regulation can be interpreted as intending to make failure sufficiently unlikely. Solvency regulation typically includes regulatory solvency capital requirements such as under Solvency II and in the SST. These can be interpreted as requiring that insurance companies demonstrate, based on a model of stochastic future developments, that they are able to “sufficiently fulfill” their insurance obligations, i.e. to produce them “in sufficiently many cases” as defined by what we call a fulfillment condition in this paper. Under Solvency II, the fulfillment condition corresponds to a 99.5% value-at-risk and, in the SST, to a 99% tail-value-at-risk (expected shortfall), in both cases over a one-year period. The production strategies that we consider thus in general provide sufficient as opposed to “full” fulfillment, with a fulfillment condition that is required to be satisfied over each one-year period, and where, in case of failure, insolvency laws trigger insolvency proceedings. [^2] In reality, regulatory solvency capital requirements and insolvency laws are not always sufficiently aligned for sufficient fulfillment to correspond to no failure in “sufficiently many cases”.
(2) Insurance companies have and use capital, for example shareholder capital, which should thus be an element of production strategies. We view capital investments as arrangements by which investors (capital providers) provide financial resources to an insurance company in exchange for promises of future benefits that depend on the outcome and cannot by themselves trigger failure. [^3] Hybrid capital, more generally, can trigger failure in specific situations. This allows capital investments to absorb risk, i.e. volatility in the insurance liabilities: if insurance liabilities turn out to be higher, benefits to capital investors can be lower. We assume that investors are prepared to enter into a capital investment if the stochastic future benefits satisfy a specific financiability condition. In line with the conception and the one-year view of Solvency II and the SST, we assume that the financiability condition is expressed as a one-year condition, and make the strong assumption that an insurance company can raise capital annually if required provided that the financiability condition is satisfied.
@@
(3) The assets of insurance companies are typically not limited to tradables, that is, “fully liquid” assets, but include also illiquid assets, for which trading is limited. We specifically consider “fully illiquid assets”, which we assume can neither be bought nor sold.444“Fully illiquid assets” in this sense are linked to the question whether insurance companies can benefit from “holding assets to maturity” (HTM) instead of having to sell them for a potentially unfavorable market price, which under Solvency II is related to “volatility adjustment” and “matching adjustment”. This implies that their use in production strategies is limited to using their cash inflows.
(4) We distinguish between production strategies that only allow for non-negative units of tradables and production strategies that also allow for negative units. The latter amounts to extending the range of available instruments by allowing to take on additional liabilities by borrowing tradables (“going short”). This requires precisely defining the additional liabilities and to consider that they can trigger insolvency. This turns out to be linked to the question whether cash inflows at future dates, specifically from fully illiquid assets and insurance contract premiums, can be used for producing insurance contract cash outflows at earlier dates. Arguably, regulatory solvency capital requirements in some cases implicitly assume that this is possible. [^5] For example, by acknowledging that the requirements account for “solvency” but not “liquidity”. As we show, this is related to whether production cost can be negative, e.g. due to premium inflows significantly exceeding cash outflows.
A key element of the framework is the precise definition of production strategies that accounts for these aspects. We assume that production strategies proceed over successive one-year periods, in line with the one-year view for example under Solvency II and in the SST. We initially consider production strategies with only non-negative units of tradables and that stop when failure occurs. Restricting to non-negative units of tradables is also worthwhile given that insurance companies often do not take on significant additional liabilities. The strategies can be extended under suitable assumptions to production strategies that can also contain negative units of tradables, and given suitable insolvency laws to production strategies that continue after failure.
One of the main results of this paper (Theorem 4.3) characterizes when the valuation according to the framework recovers the market prices of tradables, more precisely, the market values of non-negative investment strategies of tradables. This is shown for a setting that corresponds to classical no-arbitrage theory, given by production strategies that can contain negative units of tradables, no fully illiquid assets, and the “full” fulfillment condition (i.e. requiring fulfillment almost surely), but additionally allowing for capital investments. Theorem 4.3 in particular sets a requirement on the financiability condition. For more general settings, specifically allowing for only sufficient fulfillment, we would not expect that the value would be recovered.
A key idea of the paper is a two step procedure for constructing production strategies recursively backward in time over successive one-year periods. We illustrate this by a very simple concrete example. The example illustrates basic ideas of the earlier paper Möhr [8] that the framework of this paper shares but attempts to improve and generalizes. A similar setting is used in Section 5, where we consider Solvency II and the SST as special cases. For the example, consider the task of calculating the production cost $\bar{v}_{0}^{\phi}(L_{1})$ at date 00 of a production strategy $\phi$ for a given liability $L_{1}\geq 0$ at date $1$, with no cash flows between 00 and $1$. In general, $L_{1}$ could be the production cost at date $t=1$ derived in the prior backward recursion. Let $r_{0,1}$ be the risk-free rate at date 00 for a deterministic payoff at date $1$. As fulfillment condition, we assume that the liabilities $L_{1}$ need to be fulfilled by the assets $A_{1}$ with at least $99.5\%$ probability, which can be expressed by $\rho(A_{1}-L_{1})\leq 0$, where $\rho$ is the $99.5\%$ value-at-risk (Section 5).
The first of the two steps of the construction is to determine assets with value $A_{0}$ at date $t=0$ and a strategy $\phi$ until $t=1$ such that the fulfillment condition is satisfied at $t=1$. For the example, let the strategy $\phi$ consist in investing tradables with value $A_{0}\geq 0$ at date 00 risk-free until $t=1$, so that $A_{1}=(1+r_{0,1})\cdot A_{0}$. Then, if we select $A_{0}=(1+r_{0,1})^{-1}\cdot\rho(-L_{1})$, the fulfillment condition is satisfied with equality. This is because then $A_{1}=\rho(-L_{1})$ and thus using translation-invariance of $\rho$, we get $\rho(A_{1}-L_{1})=-\rho(-L_{1})+\rho(-L_{1})=0$. The second step consists in using the financiability condition to split the value $A_{0}$ into a (maximum) amount $C_{0}$ that can be raised from capital investors and the remainder, which, as we show, is the production cost $\bar{v}_{0}^{\phi}(L_{1})=A_{0}-C_{0}$ under the selected strategy $\phi$. As is explained in Section 3.1, the capital investors’ claim at date $1$ is $(A_{1}-L_{1})_{+}$. If this non-negative quantity is not positive with positive probability, no capital is used, $C_{0}=0$. Otherwise, a capital investor would likely be prepared to invest some positive amount $C_{0}>0$ at date 00 for the claim $(A_{1}-L_{1})_{+}$ at $t=1$, depending on the specific financiability condition. The production cost then is $\bar{v}_{0}^{\phi}(L_{1})=A_{0}-C_{0}$ for the following reason: given assets with value $A_{0}-C_{0}$, capital $C_{0}$ can be raised, giving assets with value $A_{0}$ that are then invested in the strategy. By construction, the financiability and the fulfillment condition then hold, so the liability $L_{1}$ can be produced.
@@
Further simplifying for illustration, let the financiability condition be given by a minimum expected excess return $\eta>0$ over the risk-free rate $r_{0,1}$, where we also ignore $()_{+}$, i.e. we have as minimum $E[A_{1}-L_{1}]=(1+r_{0,1}+\eta)\cdot C_{0}$. It then follows from a short calculation that the production cost under the strategy can be written as
$\bar{v}_{0}^{\phi}(L_{1})=\frac{E[L_{1}]}{1+r_{0,1}}+\frac{\eta}{1+r_{0,1}+\eta}\cdot\rho\left(\frac{E[L_{1}]-L_{1}}{1+r_{0,1}}\right)$ (1)
@@
This can be viewed as the sum of a “best estimate” and a “risk margin”, where the risk margin is given by the discounted capital cost for the discounted deviation of $L_{1}$ from $E[L_{1}]$. However, a separation between best estimate and risk margin is not an essential element of the framework and can in fact make calculations more complicated.
We refer to the introduction section of Engsner et al. [4] and the references provided there for some of the many papers on the valuation of (insurance) liabilities. As in [4], the current paper considers successive one-year capital requirements and insolvency and, in common also with Albrecher et al. [2], a key element is the definition and valuation of capital investments.
In the current paper, in line with Möhr [8] and unlike several other approaches to the valuation of insurance liabilities, the value of the liabilities is considered as a whole. A breakdown into e.g. a best estimate and a risk margin is not an essential element. A key difference to [8], as well as to Engsner et al. [4] and others including Albrecher et al. [2], is that the production strategies are not restricted by assuming that the assets corresponding to the capital are invested separately and specifically in a numeraire asset. The current paper is intentionally more elementary than many other papers, not assuming the existence of (“risk-neutral”) pricing measures and numeraires. A specific difference to Engsner et al. [4] is the specification of the capital investment, which in our case is set up for every one-year period with payout equal to the positive part of the value of assets over liabilities. In [4], it consists of dividend streams until the option to default is exercised by capital investors and, in our understanding, does not consider the liability cash flows after the date considered.
The paper is organized as follows. In Section 2, we set up the framework, specifically cash flows, tradables, illiquid assets, and investment strategies, with specific consideration of portfolios containing negative units of tradables. In Section 2, we set up the framework, specifically cash flows, tradables, illiquid assets, and investment strategies, with specific consideration of portfolios containing negative units of tradables. We introduce a normative condition on the market prices of tradables that is roughly comparable to “no arbitrage” and that we call “consistency”. Consistency essentially means that if, at a given point in time, the market price of a suitable strategy is not smaller than that of another suitable strategy (almost surely), then this holds at any earlier point in time.
Section 3 sets out the production of liabilities in the framework, introducing the valuation of assets and liabilities in the balance sheet, fulfillment and the financiability condition, and the central definition of a production strategy, at first limited to strategies with non-negative units of tradables and that stop whenever failure occurs. This is extended in Section 3.5 to strategies that can contain negative units and in Section 3.3 to strategies that continue after failure.
In Section 4, we analyze properties of the valuation, specifically, in Section 4.1, under which conditions the market prices of tradables are recovered. Section 4.2 looks at adding short positions to liabilities and Section 4.3 investigates production also using illiquid assets. Section 5 sets out how the valuation of insurance liabilities under Solvency II and in the SST can be seen as special cases of the framework. Section 6 concludes.